HIGH RESOLUTION LASER SPECTROSCOPY OF HAFNIUM MONOFLUORIDE.

A. G. ADAM and \underline{SCOTT HOPKINS}, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2.

High resolution laser spectra of HfF have been acquired in the visible region of the spectrum. The molecules were produced via laser ablation of a hafnium target rod, followed by reaction with SF_{6} in a pulsed supersonic jet. Several electronic transitions have been observed and analysed between 17,000 and 23,000 cm^{−1}, all yielding an Ω′′=3/2 ground state consistent with the ^{2}Δ_{3/2} ground state of HfCl ^{a}. Curiously, two electronic transitions at 19,707 cm^{−1} and 19,977 cm^{−1}, which have both been assigned as [Ω′=1/2]-X[Ω′′=3/2], exhibit ^{2}Σ - ^{2}Π structure. Work on this molecule is continuing and the results will be discussed.

aR. S. Ram, A. G. Adam, A. Tsouli, J. Levin and P. F. Bernath, J. Mol. Spectrosc. \underline{202}, 116(2000).